OpenStudy (anonymous):

Find coordinates of the point(s) of inflection for: f(x)=x^4-4x^3+6x^2 f'(x)=4x^3-12x^2+12x =4x(x^2-3x+3 Quad Form: -(-3)+-sqrt(-3^2-4(1)(3)/2(1) =3+-sqrt-3/2 Not sure what to do since the quad form. gives me an imaginary number.

OpenStudy (primeralph):

Inflection is f '' not f '

OpenStudy (anonymous):

\[f'(x)=4x^3-12x^2+12x\] \[f''(x)=12x^2-24x+12=12(x^2-2x+1)=12(x-1)^2\]

OpenStudy (anonymous):

f"=0 then sub that x value in original equation to get the y value

OpenStudy (primeralph):